A hot black body emits the energy at the rate of 16 J m^-2 s^- 1 and its most intense radiation corresponds to 20,000 Å. When the temperature of this body is further increased and its most intense radiation corresponds to 10,000 Å, then find the value of energy radiated in Jm^-2 s^-1


1
Expert's answer
2021-05-27T18:52:53-0400

We know according to Wein's displacement law,


"\lambda_m.T = b"


i.e. if "\lambda_m" becomes half temperature doubles.

Hence,


"\dfrac{\lambda_{m1}}{\lambda_{m2}} = \dfrac{20000}{10000} = \dfrac{T_2}{T_1}"


"T_2 = 2T_1"


Also, "e = \sigma T^4"


Therefore,


"\dfrac{e_1}{e_2} = (\dfrac{T_1}{T_2})^4"


"e_2 = (\dfrac{T_2}{T_1})^4e_1"


"e_2 = (\dfrac{2T_1}{T_1})^4e_1"


"e_2= 16\times 16"


"e_2 = 256Jm^{-2}s^{-1}"